Calculate net value added at market price of a firm: -

ITEMS
(Rs. IN LAKHS)
i. Sale in domestic market.
250
ii. Opening stock.
20
ii. Closing in stock
50
iii. Depreciation
15
iv. Net in direct taxes
25
v. Intermediate cost
200
vi. Export
10
1
Expert's answer
2021-10-18T18:01:14-0400

The formula for net value added at market price:

NVA at MP = (Output sold x Price Per Unit) + (Change in Stock) – (Intermediate Cost) - (Depreciation)  

where change in stock is  determined by subtracting initial stock from final stock. hence 50-20 = 30


Therefore

NVA at MP = 250 + 30 - 200 - 15

=280-200-15

=65


The answer is 65